OpenStudy (anonymous):

find the value(s) of x for which f(x)=g(x). f(x)=x^4-2x^2, g(x)=2x^2

OpenStudy (anonymous):

\[f(x)=x^2 (x^2-2)\] \[g(x)=2x^2\] when f(x)= g(x) we have f(x)-g(x)=0

OpenStudy (anonymous):

\[x^2(x^2-2-2)=0\]

OpenStudy (anonymous):

values are x=0 or x=+2, -2

OpenStudy (anonymous):

thanx